Circular No. 8467

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2005E, 2005F, 2005G, 2005H, 2005I, 2005J Three apparent supernova discoveries have been reported from unfil-

tered CCD images: SNe 2005H and 2005I by LOSS/KAIT (cf. IAUC 8465, via J. Graham and W. Li) and SN 2005J by O. Trondal and M. Schwartz (cf. IAUC 8442; Tenagra II 0.81-m telescope).

SN  2005 UT  α_{2000}  δ_{2000}  Mag.  Offset
2005H Jan. 15.17 2^{h}09^{m}38.^{s}52 −10^{o}08′43′′.6 15.9 0′′.3 E, 1′′.8 N
2005I Jan. 15.57 14 10 02.86 +17 43 22.4 18.4 21′′ W, 40′′ S
2005J Jan. 16.40 11 58 28.54 +10 01 09.2 16.5 13′′.6 E, 7′′.4 S

Additional unfiltered CCD magnitudes by the respective discoverers: SN 2005H in NGC 838, 2004 Dec. 19.16 UT, [19.0. SN 2005I in IC 983, 2004 July 2.21, [19.5. SN 2005J in NGC 4012, 2004 Jan. 21.38, [19.0; 2005 Jan. 18.42, 16.5. Note that SN 2004aq also appeared in NGC 4012.

A. Pastorello and S. Taubenberger, Max-Planck-Institut für Astrophysik (MPIA), Garching; F. Patat, European Southern Observatory; and S. Benetti, A. Harutyunyan, and N. Elias-Rosa, Osservatorio Astronomico di Padova; on behalf of the European RTN collaboration (cf. IAUC 7987), report that inspection of a spectrogram of SN 2005I, taken on Jan. 18.22 UT by M. Alises with the Calar Alto 2.2-m telescope (+ CAFOS; range 330–880 nm), shows it to be a type-II supernova, ∼ 3 months after the explosion. The spectrum is characterized by a red continuum with relatively strong and narrow P-Cyg lines of H, Ca II, Fe II, Sc II, Ba II, and Ti II. The expansion velocities, deduced by the minima of the main features (NED recession velocity of 5443 km/s assumed for the host galaxy), are 4800 km/s for Hα, 3300 km/s for Na I D, and 2500 km/s for the Fe II lines (multiplet 42). The supernova spectrum does not show clear evidence of interstellar Na lines at the rest wavelength of the host galaxy, suggesting no significant reddening. A noisy spectrum of SN 2005H is that of a young type-II supernova. The spectrum is dominated by a blue continuum with a broad Hα emission flanked by a shallow broad absorption. Hβ, Fe II 501.8and 516.9-nm, and He I 587.6-nm lines are also present with well-developed P-Cyg profiles. The expansion velocity deduced from the Hα minimum is ∼ 12000 km/s, while that deduced from Hβ and He I 587.6-nm is ∼ 10400 km/s (NED recession velocity of 3851 km/s assumed for NGC 838, a starburst galaxy).

Corrigendum. On IAUC 8465, line 4, for 2004 UT read 2005 UT

2005 January 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green